OpenStudy (anonymous):

2. A system of equations is shown below. {5x+ 9y=-54 {2x+ y=20 a.) Find the solution to the system of equations using either substitution or the linear combination method. Show all your work.

OpenStudy (texaschic101):

2x + y = 20 --- is that what this equation is ??

OpenStudy (anonymous):

yes, it is all in one bracket but i didn't know how to do that

OpenStudy (texaschic101):

2x + y = 20 y = -2x + 20 understand so far ?

OpenStudy (texaschic101):

now sub -2x + 20 in for y in the other equation 5x + 9y = -54 5x + 9(-2x + 20) = -54 -- distribute through the parenthesis 5x - 18x + 180 = -54 -- combine like terms -13x + 180 = -54 -- subtract 180 from both sides -13x = -54 - 180 -13x = - 234 x = 18 now sub 18 in for x in either of the equations 2x + y = 20 2(18) + y = 20 36 + y = 20 y = -36 + 20 y = -16 check.. 2x + y = 20 2(18) -16 = 20 36 - 16 = 20 20 = 20 (correct) solution is : x = 18 and y = -16 any questions ?

OpenStudy (anonymous):

no ma'am thank you so much again! i need help with one more, can you help? c:

OpenStudy (texaschic101):

I can try